# An Analysis of Hyper-Parameter Optimization Methods for Retrieval Augmented Generation

### v1

@@ 6 Analysis
Figure 4 depicts per-dataset maximal answer correctness scores for each generative model (maximum of all $162/3=54$ configurations the model participates in). Interestingly, the best RAG configuration for all datasets involves Llama when optimizing by LLMaaJ-AC and Granite or Mistral when optimizing by Lexical-AC. This difference stems from the metrics, as Lexical-AC is recall-oriented while LLMaaJ-AC accounts for both precision and recall. This difference highlights the importance of selecting an objective metric that correctly reflects use case preferences, as optimal configuration may vary for different objective metrics.
@@
Figure 5 depicts the relationship between context correctness and answer correctness. The two measures are correlated, and specifically, when retrieval quality is low, so is answer quality. Also evident is the great influence of generative model choice on answer quality.
@@
We also track the cost of each HPO algorithm, in terms of number of embedded tokens and number of input and output tokens during generation. These counts are accumulated per-algorithm over all explored configurations, where indices that are reused in multiple explored configurations are accounted for only once.
The number of different indices created by each algorithm is the most significant factor in embedding cost. TPE and Random behave the same, as they have no preference to reuse existing indices nor vice versa. Greedy-R and Greedy-R-CC behave the same: at start, they explore only different indices, hence they are more expensive, but once the index configuration is fixed, they become cheaper. Greedy-M starts with one index and iterates first over generation models, so it is initally less costly, but over the next iterations it behaves similarly to TPE and Random overtaking Greedy-R and Greedy-R-CC. These behaviors are consistent across all datasets. We do not find a significant difference in generation token cost across algorithms. Results for both metrics are in Appendix C.
@@ 7 Conclusion
Our work presented a comprehensive study of multiple RAG HPO algorithms, exploring their performance over datasets from various domains, using different metrics. The results can be summarized into important guidelines for practitioners.
We demonstrated the importance of RAG HPO, showing that it can boost performance significantly compared to an arbitrary starting point (by as much as $20$%). As this holds for a simple RAG pipeline, the effect on a more complex (or agentic) RAG pipelines with many more hyper-parameters is likely to be larger.
We emphasize the need for an automated approach. For example, the impact of the optimization metric on the optimal RAG configuration (see Figure 4) dictates that when changing metrics even the same RAG pipeline must be re-adjusted to the new metric with HPO. Similarly, changes in available models or evolving datasets require revising the chosen RAG configuration.
@@
We take initial steps in improving efficiency of RAG HPO, evaluating development set sampling, and plan to further explore this in future work. We also contribute a new RAG dataset, of enterprise product documentation, for use by the community.
8 Limitations
@@
Our experiments optimized a simple linear RAG workflow. Additional RAG approaches may introduce additional hyper-parameters to the mix, such as filters, rerankers, answer reviewers, HyDE etc. Moreover, our prompts were fixed per generation model and we did not explore prompt optimization.
@@ Appendix B Additional results
Figure 6 details development sampling results for the Lexical-AC metric.
@@ Appendix C Embedding and Generation Costs
Figure 7 details accumulated numbers of (a) embedded tokens and (b) number of tokens used in generation part for HPO algorithms overall tested configurations.
@@
(a) Total number of tokens for chunks sent to the embedding models.

### v2

@@ 6 Analysis
Figure 4 depicts per-dataset maximal answer correctness scores for each generative model (maximum of all $162/3=54$ configurations the model participates in). Interestingly, the best RAG configuration for all datasets involves Llama when optimizing by LLMaaJ-AC, and Granite or Mistral when optimizing by Lexical-AC. This difference stems from the metrics, as Lexical-AC is recall-oriented while LLMaaJ-AC accounts for both precision and recall. This highlights the importance of selecting a target metric that correctly reflects the use case preferences, as optimal configurations may vary for different target metrics.
@@
Figure 5 depicts the relationship between context correctness and answer correctness. The two measures are correlated – when retrieval quality is low, so is answer quality. Also evident is the influence of the generative model choice on answer quality.
@@
The cost of each HPO algorithm was tracked by counting the number of tokens embedded when creating each index, and the number of input and output tokens used in generation. These counts are accumulated per-algorithm over all explored configurations till some specific iteration. Indices that are reused in multiple explored configurations are accounted for once. Detailed plots of these counts are in Appendix C.
Overall, the generation cost was about the same for all algorithms and datasets. The embedding cost, being a function of the number of different indices created by each algorithm, behaved similarly across datasets with some variations between algorithms. In this regard, both TPE and Random have no built-in preference to reuse already constructed indices. In contrast, both Greedy-R and Greedy-R-CC create a new index in each iteration till they have optimized all index-related parameters, and are therefore initially costly. Greedy-M is initially cheap, since starts by optimizing the generative model parameter using just one index. It is therefore preferable if cost is crucial and the limit on the number of explored configuration is low.
@@ 7 Conclusion
Our work presented a comprehensive study of multiple RAG HPO algorithms, exploring their performance over datasets from various domains and using multiple metrics, and demonstrated that RAG HPO can boost performance significantly. Compared to an arbitrary choice of RAG configurations, $10$ HPO iterations improve performance by as much as $20$% (see Figure 2, comparing performance at iterations $1$ and $10$). Since this holds for simple RAG pipelines, the effect on more complex (or agentic) RAG pipelines with many more hyper-parameters is likely to be even greater, a subject for further work.
@@
Our results highlight the importance of the chosen optimization objective, as different choices of metrics correspond to different optimal RAG configurations. Jointly optimizing multiple metrics is an interesting direction for future work.
We take initial steps in improving the efficiency of RAG HPO, by sampling the development set, and plan to further explore this in future work. We also contribute a new RAG dataset, of enterprise product documentation, for use by the community.
Limitations
@@
Our experiments optimized a simple linear RAG workflow comprising retrieval followed by generation. Additional RAG approaches may introduce additional hyper-parameters to the mix, such as filters, rerankers, answer reviewers, HyDE etc. Moreover, our prompts were fixed per generation model and we did not explore prompt optimization.
@@ Appendix A Datasets
A.1 WatsonxQA additional details
As stated in the body of the paper, the WatsonxQA benchmark includes $75$ QA pairs and gold document labels, of which $50$ were generated synthetically. The benchmark contains five fields: a question, a gold answer and, for computing context relevance metrics, the gold passage id and its content. Some example benchmark entries are given in Figure 12. Synthetic generation was operated using falcon-180b model, and then manually filtered and reviewed for quality - the methodology we used is detailed in Yehudai et al. (2024).
The prompt used for the synthetic generation is detailed in  Figure 13.
@@ Appendix B Additional results
Figure 6 details test results for the Lexical-FF metric. Figure 7 details development sampling results for the Lexical-AC metric. See §5 for a discussion of the main results.
Figure 6: Per-iteration performance of all HPO algorithms on the test sets of five datasets, optimizing answer faithfulness. The dashed black lines denote the best achievable performance on each test set. See §5 for a discussion of the main results.
@@ Appendix C Embedding and Generation Costs
Figure 8 details accumulated numbers of (a) embedded tokens and (b) number of tokens used in generation part for HPO algorithms overall tested configurations. See §6 for a discussion of costs.
@@
(a) Total number of tokens from chunks sent to the embedding models.
@@ Appendix F Generation Details
Question: What are the natural language processing tasks supported in the Product library?
Gold answer: The Product library supports the following natural language processing tasks: language detection, syntax analysis, noun phrase extraction, keyword extraction and ranking, entity extraction, sentiment classification, and tone classification.
Gold context-id: EN-001
Gold passage: The following natural language processing tasks are supported as blocks or workflows in the Product library:
* [Language detection]
* [Noun phrase extraction]
* [Keyword extraction and ranking]
* [Entity extraction]
* [Sentiment classification]
* [Tone classification]
  Question: What happens to unsaved prompt text within Product, and how long does it persist on the webpage before being deleted?
Gold answer: The prompt text remains unsaved unless the user decides to save their progress. While unsaved, the prompt text persists on the webpage until a page refresh occurs, upon which the text is automatically deleted.
Gold context-id: EN-101
Gold passage: Privacy of text in Product during a session
Text that you submit by clicking Generate from the prompt editor in Product is reformatted as tokens, and then submitted to the foundation model you choose. The submitted message is encrypted in transit.
Your prompt text is not saved unless you choose to save your work.
Unsaved prompt text is kept in the web page until the page is refreshed, at which time the prompt text is deleted.
Figure 12: Examples of benchmark entries in WatsonxQA.
Given the next [document], create a [question] and [answer] pair that are grounded in the main point of the document, don’t add any additional information that is not in the document. The [question] is by an information-seeking User and the [answer] is provided by a helping AI Agent. [document]: [A WatsonxQA document example] # Response: [question]: What is a token limit? [answer]: Every model has an upper limit to the number of tokens in the input prompt plus the number of tokens in the generated output from the model (sometimes called context window length, context window, context length, or maximum sequence length.) [document]:
Figure 13: The prompt used for synthetic data generation. The prompt consists of instruction followed by $K=3$ examples of document and generated QA.